OpenStudy (anonymous):

solve for v |3v+18|=0

OpenStudy (anonymous):

You make two equations, |3(-v) +18| = 0 and |3v + 18| = 0

OpenStudy (anonymous):

Oops! The first one is |-3v + 18| = 0, sorry.

OpenStudy (anonymous):

Now we just get the "v's" by them selves. Lets solve the first equation.

OpenStudy (anonymous):

|-3v + 18| = 0 -3v + 18 = 0 -3v = -18 v = 6

OpenStudy (anonymous):

Can you do the second one? Tell me if you need help.